OpenStudy (anonymous):

what is the vertex form of y=x^2-6x+6? Answers will be in comments.

OpenStudy (anonymous):

a) y=(x-3)^2-15 b)y=(x-3)^2-9 c) y=(x-3)^2-6 d)(x-3)^2-3

OpenStudy (anonymous):

@skinny23